# The universal family of punctured Riemann surfaces is Stein

### v2

@@ Abstract.
We show that for any pair of integers $n\geq 1$ and $g\geq 0$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
@@ 1. Introduction
(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$.
In this paper we prove the following result.
Theorem 1.1.
@@
The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [22, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [21]). Note also that $V(g,0)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,0)$.
The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1.1), corresponding to the punctures of $V(g,n)$, are called canonical sections in the literrature. The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [19] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$. See also [20]. Earle and Kra proved in [7, p. 50] that for any $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
Theorem 1.1 has interesting consequences for function theory and complex geometry. By classical results of Remmert, Bishop, and Narasiman (see [9, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$ (which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$). When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ by the optimal embedding theorems of Eliashberg and Gromov [8] and Schürmann [24]. Since every fibre of $V(g,n)$ is a noncompact Riemann surface of finite conformal type, and hence is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
Problem 1.2.
@@
Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$. The following is an immediate corollary to this observation, Theorem 1.1, and the main result of Oka theory [9, Theorem 5.4.4]. For Oka manifolds, see [9, Chap. 5] and the surveys [10, 12, 13].
Corollary 1.3.
@@
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [9, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
@@
In this connection, we mention that in [11] the Oka principle was established for families of maps from very general families of open Riemann surfaces $(M,J_{t})$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space, which is not a complex manifold in general. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [11, Theorem 1.6] are merely continuous or smooth in the parameter. In [14] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds. Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$. This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
2. Proof of Theorem 1.1
@@
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ in (1.1), Theorem 1.1 is an immediate consequence of the following more general result with arbitrary Stein manifold as the base.
Theorem 2.1.
@@
Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
@@
Proof.
Let $\pi:Z\to X$ be as in the theorem. We may assume that $X$ is connected. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
Recall the following result (see Grauert and Remmert [16, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [16, Chap. A.3.5]). Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
By a theorem of Siu [25], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [16, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [15] will then imply that $Z\setminus H$ is Stein.
@@
(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
@@
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [9, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
@@
This completes the proof of Theorem 2.1, and it also proves Theorem 1.1. ∎
A minor modification of the proof of Theorem 2.1 gives the following more general result.
Theorem 2.2.
Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. Then, the domain $\Omega=Z\setminus H$ is Stein.
Sketch of proof.
The restricted projection $\pi:H\to X$ is a proper holomorphic map. Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover). The proper mapping theorem of Remmert [23] (see also Chirka [6, p. 29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [16, Theorem 1 (d), p. 125]. By Siu’s theorem [25], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$. By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [16, Theorem 5, p. 129]. The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave further details to the reader. ∎
Acknowledgements. Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia. I wish to thank Finnur Lárusson for posing the question answered by Theorem 1.1 in a private communication.

### v3

@@ Abstract.
We show that for any pair of integers $g\geq 0$ and $n\geq 1$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold. We describe its basic function theoretic properties and pose several challenging questions.
@@
(1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$. If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [28, Theorem 5.4.3]). In this paper we prove the following result.
Theorem 1.
@@
This is a special case of Theorems 8 and 9 given in the sequel. The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [28, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [27]). The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1), corresponding to the punctures of $V(g,n)$, are called canonical sections. The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [25] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$. See also [26]. Earle and Kra proved in [10, p. 50] that for $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
Theorem 1 has interesting consequences for function theory and complex geometry. By classical results of Remmert, Bishop, and Narasiman (see [15, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$. When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [11] and Schürmann [31], or the exposition in [15, Secs. 9.3–9.4]). Since every fibre of $V(g,n)$ is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
Problem 2.
@@
Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$. The following is an immediate corollary to this observation, Theorem 1, and the main result of Oka theory [15, Theorem 5.4.4]. For Oka manifolds, see [15, Chap. 5] and the surveys [13, 12, 17].
Corollary 3.
@@
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [15, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial. Since $V(g,n)$ is Stein, the Oka–Grauert principle ([19] or [15, Theorem 3.2.1]) implies the following.
Proposition 5.
Every holomorphic vector bundle on $V(g,n)$ for $n\geq 1$ is holomorphically trivial.
Corollary 6.
Assume that $g\geq 0$ and $n\geq 1$.
  1. (a)
    There exists a nowhere vanishing holomorphic vector field $\xi$ on $V(g,n)$ which is tangent to the fibres of the projection $\pi:V(g,n)\to T(g,n)$, that is, $d\pi(\xi)=0$.
  2. (b)
    With $\xi$ as in (a), there exists a holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$. In particular, $\theta$ is nowhere vanishing on the tangent bundle to any fibre of $\pi$.
Proof.
Part (a) follows by applying Proposition 5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$. To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
$0\longrightarrow\ker d\pi\lhook\joinrel\relbar\joinrel\rightarrow TV(g,n)\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}H:=TV(g,n)/\ker\pi\longrightarrow 0.$
Here, $TV(g,n)$ denote the tangent bundle of $V(g,n)$. By Cartan’s Theorem B the sequence splits, i.e. there is a holomorphic vector bundle injection $\sigma:H\hookrightarrow TV(g,n)$ such that $\alpha\circ\sigma=\mathrm{Id}_{H}$. Hence, $TV(g,n)=\ker d\pi\oplus\sigma(H)=\mathbb{C}\xi\oplus\sigma(H)$, where $\xi$ is as in part (a). The unique holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$ and $\xi=0$ on $\sigma(H)$ clearly satisfies part (b). ∎
By the Gunning–Narasimhan theorem [23], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 6 (b), the following is a natural question.
Problem 7.
Let $n\geq 1$. Does there exist a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is an immersion?
By [14, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points. The problem is to find a function $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$. Note that [16, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is a holomorphic immersion. Problem 7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$. However, this is not the only problem. Since the submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$, there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
@@
In [16] the Oka principle was established for families of maps from very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [16, Theorem 1.6] are merely continuous or smooth in the parameter. In [18] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds. Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$. This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
Another interesting question is whether the Riemann surfaces $M_{t}$ in the universal Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$. For background, see [29] and [4]. By [16, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter. It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature. Each single surface in the family can be made such by [6].
@@
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$, Theorem 1 is an immediate consequence of the following result with arbitrary Stein manifold as the base. See also Theorem 9 for a more general result.
Theorem 8.
@@
Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
@@
Proof of Theorem 8.
Let $\pi:Z\to X$ be as in the theorem. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
Recall the following result (see Grauert and Remmert [21, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [21, Chap. A.3.5]). Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
By a theorem of Siu [32], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [21, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [20] will then imply that $Z\setminus H$ is Stein.
@@
(2) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
@@
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [15, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
@@
This completes the proof of Theorem 8, and it also proves Theorem 1. ∎
A minor modification of the proof of Theorem 8 gives the following more general result.
Theorem 9.
Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. Then, the domain $\Omega=Z\setminus H$ is Stein.
Proof.
The restricted projection $\pi:H\to X$ is a proper holomorphic map whose fibres are compact Riemann surfaces. Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover). The proper mapping theorem of Remmert [30] (see also Chirka [9, p. 29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [21, Theorem 1 (d), p. 125]. By Siu’s theorem [32], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$. By the argument in the proof of Theorem 8, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [21, Theorem 5, p. 129]. The proof can now be completed by a similar argument as in the proof of Theorem 8, and we leave further details to the reader. ∎
Acknowledgements. Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia. I wish to thank Finnur Lárusson for posing the question answered by Theorem 1 in a private communication.